OpenStudy (yanasidlinskiy):

Can someone help me?? Find the geometric mean of the pair of numbers. 242 and 8 A. 44 B. 49 C. 1872 D. 54

ganeshie8 (ganeshie8):

multiply the numbers, then take square root

ganeshie8 (ganeshie8):

geometric mean of \(a\) and \(b\) = \(\large \sqrt{a\times b}\)

OpenStudy (yanasidlinskiy):

How would I know if it's a or b??Like...would 242 be a or b??

ganeshie8 (ganeshie8):

that's a good question :) the answer is it doesn't matter cuz \(a\times b = b\times a\)

OpenStudy (paki):

yeah geometric mean is the square root of their mean....

ganeshie8 (ganeshie8):

order doesn't matter when you multiply...

OpenStudy (yanasidlinskiy):

Let me see what I get.. \[\sqrt{242*8}\]

OpenStudy (yanasidlinskiy):

Is that what it's going to look like??

ganeshie8 (ganeshie8):

Yep !

OpenStudy (yanasidlinskiy):

Ok. So now I have: \[\sqrt{1936}\] What do I do next?

ganeshie8 (ganeshie8):

simplify : \(\large \sqrt{242 \times 8 } = \sqrt{121 \times 2 \times 8} = \sqrt{121\times 16} = 11\times 4 = 44\)

ganeshie8 (ganeshie8):

plugging into calculator works too^

OpenStudy (yanasidlinskiy):

Wait..What was I suppose to do? What did you do?? Yea..Thanx about the caluclator..It's on my bed though...I use my *online*.

OpenStudy (yanasidlinskiy):

Where did you get the 2 and 121 from?

OpenStudy (yanasidlinskiy):

Anyways...Thanx sooooo much for helping me!!!:) I appreciate your help!!:)

ganeshie8 (ganeshie8):

Hey sorry I am back

ganeshie8 (ganeshie8):

\(242 = 121\times 2\)

OpenStudy (yanasidlinskiy):

Thanx!!:) Again. I thought you were never going to reply back!:) Thanx for replying and showing me where you got that from!!:) It's ok!:) I forgive ya:)

ganeshie8 (ganeshie8):

lol thank you :)

OpenStudy (yanasidlinskiy):

Your welcome:0 I'll post a new one..

ganeshie8 (ganeshie8):

Okay sure...